OpenStudy (anonymous):

Hey anyone can help me this problem: the sum of three numbers is 90. find the numbers if the second number is 2 more than twice the first and the third is two less than the sum of the first two.

OpenStudy (anonymous):

x+y+z=90 ............................eq 1 2x+2=y i.e 2x-y=-2................eq 2 x+y-2=z i.e. x+y-z=2 ...............eq 3 eq 1 - eq 3 implies 2z = 88 i.e z =44 eq 3 becomes x +y -44 = 2 i.e x+y = 46.......... eq 4 eq 2 + eq 4 implies 3x = 44 so x = 44/3 put in eq 2 88/3 - y = -2 so y = 88/3 + 2 or y =94/3